OpenStudy (anonymous):

simplify leaving answer in surd form ... radical 125 / 5

OpenStudy (ash2326):

is it \(\sqrt(125/5)\)??

OpenStudy (anonymous):

√125 divided by 5

OpenStudy (ash2326):

ok 125=5*5*5 so \(\sqrt{125}\)=\(5\sqrt5\) \(\sqrt{125}/5=\sqrt{5}\)